### v1

# Language models should be subject to repeatable, open, domain-contextualized hallucination benchmarking

Abstract
Plausible, but inaccurate, tokens in model-generated text are widely believed to be pervasive and problematic for the responsible adoption of language models. Despite this concern, there is little scientific work that attempts to measure the prevalence of language model hallucination in a comprehensive way. In this paper, we argue that language models should be evaluated using repeatable, open, and domain-contextualized hallucination benchmarking. We present a taxonomy of hallucinations alongside a case study that demonstrates that when experts are absent from the early stages of data creation, the resulting hallucination metrics lack validity and practical utility.
1 Introduction
Language models should be subject to repeatable, open, domain-contextualized hallucination benchmarking. The current environment of hallucination benchmarking is hampering the development of models that could perform better and with fewer errors and hallucinations. Specifically, because benchmarks are (a) not open; (b) not repeatable; (c) are either domain ignorant or ablate expertize behind private walls, it is not possible for researchers and commercial developers to openly compare model performance in the pursuit of better functioning, safer, and less-carbon intensive frameworks. By developing open and repeatable benchmarks, the language modeling community can establish a shared set of tasks that support model training and experimentation prior to submitting models to public evaluation. When these benchmarks incorporate expert knowledge, they help ensure that resulting models are both valid and practically useful.
To identify and mitigate hallucinations in contemporary LLMs, a crucial first step is the development of a process to create a golden dataset that supports domain-specific hallucination benchmarks. One key lesson we learned—the hard way—is that domain expertise must be present from the inception of the dataset. Without domain expert involvement in the early stages of data creation, hallucination metrics lack scientific validity and real-world utility.
This paper proceeds as follows: we begin by outlining the problem space of hallucination benchmarking, summarizing the state of the field, providing a working definition of model hallucination, and providing three examples the consequences of model hallucination in the real-world. We then argue why a repeatable, open, domain-contextualized approach will produce better outcomes. We then demonstrate how expert responses differ from model-generated text, drawing on one case study from medicine and a second from politics. We conclude by evaluating alternative perspectives.
2 Language Model Hallucination
Although there are several descriptive features about language models — the amount of time or carbon to train a model, the speed with which a model generates tokens at inference, inter alia — we focus specifically on controlling incorrect or misleading text, i.e., hallucination benchmarking. [^1] See Appendix B for a brief survey of model performance and accuracy. Recent advances in open and repeatable benchmarks (e.g., srivastava2023, ; bbeh2025, ) have helped to constrain general-purpose, model-generated text. The result is that LLM-generated text has become increasingly, though only seemingly, coherent. But the shine, has come off the stochastic parrot — simply generating seemingly coherent language is no longer sufficient, and language models are being pushed to produce outputs that serve concrete business, policy, medical and scientific purposes. Stakeholders increasingly want language models to take expert-like actions in high-stakes domains, but current models are not ready to operate at that level of rigor or reliability. Indeed, as models grow larger, the hallucination problem appears to be worsening (Zhou2024larger_and_more, ).
Early efforts in LLM hallucination benchmarking primarily centered on summarization tasks that compared model-generated summaries against reference texts (e.g., CNN and DailyMail (hermann2015, ; narayan2018, ) and XSum (narayan2018, )) to assess factual consistency. As LLMs evolved and were used to generate seemingly novel text, new benchmarks that did not rely upon scoring against reference texts became necessary. When benchmarking metrics scored only perceived coherence or helpfulness, teams’ trained models to score better on these measures, but with the side effect of wild hallucinations — entirely ungrounded references, links, and factual claims (see, for example, subsection 2.2) (huang2025survey, ). Recent advancements have moved from evaluating the perceived coherence of language, to assessing the factual accuracy of model-generated text in zero-shot question answering (see, e.g., Hughes_Vectara_Hallucination_Leaderboard_2023, ; li2023haluevallargescalehallucinationevaluation, ).
We highlight two recent advancements in hallucination detection: the Vectara hallucination leaderboard and HaluEval. The Vectara hallucination leaderboard provides an open and continuously updated platform for evaluating hallucination rates across LLMs using document summary metric, the Hughes Hallucination Evaluation Model (HHEM) (Hughes_Vectara_Hallucination_Leaderboard_2023, ; hhem-2.1-open, ). HHEM assesses factuality using a zero-shot question-answering setup over grounded documents; humans define ground-truth, binary judgments of correctness. In a separate project, HaluEval collects 45,000 human-annotated samples across multiple generation tasks (e.g., question answering, summarization, open-domain generation), distinguishing between extractive, abstractive, and factual hallucinations (li2023haluevallargescalehallucinationevaluation, ). Crucially, HaluEval 2.0, provides a domain-relevant benchmarks across biomedicine, finance, science, education, and as well as an open category (li2024dawndarkempiricalstudy, ).
Despite these advancements, we observe that the field has focused on only a small subset of possible dimensions from the full range available to evaluate and constrain models. These selected dimensions often include tasks like causal reasoning, geometric rotation, and calendar scheduling under conditions of incomplete information (see, e.g. bbeh2025, ). Certainly these are difficult tasks for machines — without the ability to intervene, data simply cannot possess the full compliment of counterfactual measurements to define a tractable loss function — but they are trivial for any human. It is unsurprising that model performance tends to degrade on zero-shot tasks that fall outside the scope of these evaluations. The use of ever larger models has resulted in improved performance on in-benchmark tasks but also increased hallucination on out-of-benchmark tasks (Zhou2024larger_and_more, ).
2.1 Defining LLM Hallucination
While there is a considerable body of work surrounding language model hallucination, the field has not settled on a single operational definition of what, specifically constitutes a hallucination (see e.g., ziwei2023, ; huang2025survey, , for a survey of the field). The term, hallucination, arises in 2017 in the context of a data-to-text generation task (wiseman2017challenges, ), the concept features in abstractive summarization in 2020 (maynez2020faithfulness, ), and features prominently in Stochastic Parrots (bender2021dangers, ). To create a point of reference within this paper, we define language model hallucinations in the following way:
Definition 1
An LLM hallucination is a generated text statement presented without qualifying language that is: incoherent, factually incorrect, or unverified in the judgment of a human expert capable of making that assessment.
Qualifying Language
A response that does not include qualifying language fails to convey the probabilistic nature of the output, instead presenting ideas as definitive facts. Responses that appropriately include qualifying language — for example, in medicine — might use phrases such as “Those symptoms could be a sign of [diagnosis] but they could also be signs of [alternative diagnosis],” or “It is important to consult with a healthcare provider…” to communicate uncertainty and encourage additional verification.
Incoherence
The greatest trick language models ever pulled was convincing the world they were coherent. Current generation language models produce syntactically correct, seemingly coherent strings of text (bender2021dangers, ). In particular, the challenge this presents is that “…Our perception of natural language text, regardless of how it was generated, is mediated by our own linguistic competence and our predisposition to interpret communicative acts as conveying coherent meaning and intent” (bender2021dangers, , p. 616). [^2] The authors have first-hand language of this type of incoherent language: when checking results of a statistics test they administered, they asked an LLM to generate solutions for all the test questions. In one case, the answers relied on theorems that were not presented in the course; in another case, the language models produced text that engaged in a loop of reasoning without advancing the solution concept toward a conclusive statement. In other words, because the text sounds like language that a human produces, we interpret the text to have meaning and intent.
Factually Incorrect
LLM text that misstates a verifiable fact is perhaps the most obvious form of hallucination. Huang et. al. identity two types of fact-based hallucinations: (1) factual contradiction where text presents statements that can be verified with real-world information; and (2) factual fabrication where models make up facts to support statements, but these facts are inconsistent with verifiable real-world knowledge. In subsection 2.2 we develop extended examples of this type of factual fabrication. Presently, factual contradiction is undertaken by selecting only “high-quality” data sources (see Lucy et. al. for important context surrounding data filtering (lucy2024aboutme, )), Retrieval Augmented Generation (RAG), few-shot training and knowledge graph based approaches, and, recently agent-based approaches. The profusion of approaches belies the reality that this type of factual incorrectness is the easiest type of problem to identify, and so has been the locus of considerable development.
Unverified
Unverified responses that are those that either fail to cite a source, or cite an external source that is insufficient to corroborate the claim — whether to confirm or contradict it (ziwei2023, ) (huang2025survey, ). For example, if a user asks about eligibility requirements to vote in California, an LLM hallucination may either fail to cite a source altogether or cite one — such as a nonprofit organization or the California Secretary of State — where the source cited is insufficient to verify the validity of the LLM response.
2.2 Consequences of Language Model Hallucination
Language model hallucination is not simply a problem to be solved so that models score better on benchmarks. These models power widely used consumer products, and when users read and act on hallucinated output, there is a real risk of harmful consequences. Because communication is so core to the human experience, syntactically correct language serves as a credibility hack (chomsky1976, ; christiansen2008, ), effectively jailbreaking our perceptive system to imbue what seems like coherence from a stochastic parrot (bender2021dangers, ). LLM models can hide their lack of coherence within the written word — LLMs produce text that impersonates the writing of experts, and are imbued with “expertise” by readers. When combined with human decision making which is well-known to be prone to using lower-cost heuristic evaluation (simon1955, ; simon1956, ; kahneman2011, ) — even among high-stakes decision makers with very strong financial and political incentives for decision accuracy (hafner2013, ) — this jailbreak creates a profound risk for misinterpretation, and evaluation of model-generated language (bender2021dangers, ). In this section, we introduce three examples, two of which we present to domain experts.
Voters Seeking Political Process Information
The manner for how elections are conducted is constitutionally devolved to the states (Article 1, Section 4, US Constitution). Although federal law stipulates an common election day, states are largely responsible for determining the means and methods for how elections are run. Consequently, if a voter were to move from one state to another, it is possible that the election process might look quite different. The differences may include: which state agency handles registration; registration requirements and dates; availability of absentee or early voting; voter identification requirements; among others. In this context, it would be quite reasonable for a voter to query an LLM with a question like, “What documentation do I need to vote, and when does early voting start in [my location]?” Queries of this type generated documented instances of hallucination that would lead people to the wrong early voting date. These hallucinations persisted even when the model was instructed to describe the “click-path” that was used to navigate from an official government source to the specific resource used to produce an election date; the model generated text that traversed to a non-existent series of sites (hughes2025, ). For some period of time ChatGPT 3.5-x returned a caution to these political information queries, encouraging users to confirm information with their local elections agencies; however, these cautions are no longer present for users who are using ChatGPT 4o in April 2025.333OpenAI is also aware of the limitations of ChatGPT for expert users in the political space. In 2024, the company prohibited politicians and lobbyists from using ChatGPT for official campaign activities(NPR2024, ). There is also evidence that LLMs generated inaccurate and misleading responses for voters during the 2024 U.S. presidential primary (nyt2025, ).
Medical Diagnosis
Medical diagnosis is a challenging task. Patients noisily introspect their own state, commonly exhibit social desirability bias, and frequently present with comorbid diseases that are caused by a complex relationship between genetics, lifestyle, and proximate causes. Front line clinicians — nurse practitioners and general practitioners — develop expertise that incorporates medical science and real-world clinical training and use this expertise to ask winnowing questions and to order informative tests to arrive at an manageable set of possible diagnoses. With this set, doctors propose medical interventions and observe how patients respond.
Compared to clinicians, language models are missing huge swaths of the diagnostic toolkit. Models do not have access to long clinical histories or family or social context. They do not have the ability to manipulate, either physically to check the severity of an ankle sprain, or through prescription to see whether a fever responds to acetaminophen. And, models cannot order labs or imaging to distinguish between hidden states that are unknowable to a patient. [^4] In spite of this, some may may use LLMs as complete substitutes for visits to the doctor’s office. The costs of asking language model are substantially lower than attending a clinic: language models do not activate the same stigma responses as human-to-human conversation; querying a language model is substantially easier and quicker than scheduling and traveling to an in-person clinical visit; and, for under-resourced populations with limited clinical access, language models may not be an approximation or substitute for clinical care, it my be the only option. It is little surprise, then, that model-based diagnostic accuracy is substantially lower than diagnostic accuracy of expert doctors (takita2025systematic, ). To the extent that existing work includes medical expertise in the creation and benchmarking of language models, they are largely “bolting-on” doctors’ expert review at the end of the diagnostic pipeline (see, e.g., wu2024llmsciterelevantmedical, ; kim2025, ).
AI in the Criminal Justice System
Technology has a long history of being used in the field of criminal justice. Before the current era of generative artificial intelligence, AI/ML tools were mostly relegated to tasks such as surveillance, event prediction and crime prevention. When applied to such use cases, these tools were repeatedly found to reinforce or enable new systemic biases, compromise the privacy rights of citizens, and in many cases simply proved to be ineffective at their stated purpose (buolamwini2018gender, ; raji2020saving, ; taboh2021clearview, ). A shift toward more advanced models has taken place in recent years. For example, image restoration models can be used on biometric recognition (norman2024investigation, ), and LLMs (like ChatGPT) are increasigly integrated into lawyers’ workflows (tan2023chatgpt, ; izzidien2024llm, ). The efficacy of both applications is limited. For example, in 2025, lawyers who used an LLM tool to generate legal briefs were disciplined in court after it was discovered that documentation submitted contained hallucinated case law (lawhallucination, ). The potential for scaled harm is clear: in cases where the hallucinated citation is not detected, the error is carried forward until it is discovered, setting the foundation for potentially catastrophic and cascading legal crisis.
3 Solution Space
Here, we present the requirements of a scientific solution to address model hallucination. We build from the basic requirement of internal evaluation using repeatable tests; to the open sharing of tests and test performance to enable replication between research teams; and finally to our core argument that benchmarks need to contain carefully solicited expertise from experts in the domains where the model will be deployed. We present arguments for each of these three objectives separately, but would like to be quite clear that none of these are sufficient to serve as a whole benchmark. Instead, each is necessary to produce a functioning benchmark, and only when all three are objectives are implemented in concert can a benchmarking system effectively guide model development.
Repeatable
Language model developers should have repeatable benchmarks that allow for internal experimentation and evaluation prior to submitting models for public benchmarking. Without repeatable internal benchmarks, teams must repeatedly submit against high-visibility, public evaluation benchmarks (e.g., Hughes_Vectara_Hallucination_Leaderboard_2023, ; li2023haluevallargescalehallucinationevaluation, ; hong2024hallucinationsleaderboardopen, ; bbeh2025, ) creating perverse consequences for the advancement of the science of language models.
When consumers choose model adoption based on highly-visible leader-boards, scientific and profit incentives are muddied. Recall, for example, that in 2015 a research team claimed a 10.2% relative improvement over the state of the art. As would come to light through several revisions to the paper and considerable public hand-wringing by the contest organizers (ilsvrc2015announcement, ), the research team had submitted 200 distinct models to the scoring dataset, more than all other entrants combined (simonite2015, ). If development teams do not have repeatable benchmarks that can be used while building and training models, and instead have to evaluate their work on public scoring benchmarks, this problem will assuredly continue to arise (wiggers2025crowdsourced, ). Accelerating this problem are additional selection pressures: developers want to work on competitive teams, and technical decision makers want to purchase an implement the best performing models.
Open
Openness and repeatability are principal tenants of scientific investigation. As a field, computer science has a long history of building computing tools in the pursuit of this openness (e.g., knuth1984texbook, ). Not only does openness allow for the development and transmission of ideas between researchers, research groups and universities, but it also allows for open evaluation and repeatable execution of tests conducted against data (kluyver2016jupyter, ). The open science movement that makes compute environments, scientific code, and data widely available has become de rigueur for publication across many scientific communities, speeding discovery and error correction, and reducing malfeasance (christensen2019, ). Benchmarking openness, as we intend it, combines aspects of scientific openness and open source software (OSS).
Consistent with scientific openness, benchmark openness should be open about the benchmark questions, the benchmark code, and the benchmark scores. This type of scientific openness supports the repeatable benchmarking position presented in the last subsection. When consumers are exposed to technology products that are mere days off of the scientific development floor, these products must necessarily contain aspects of scientific evaluation. Openness about how a model performs against different benchmarks is a minimum “safety evaluation” that is due to consumers who are going to use a product. Just as medication labels contain statements of side effects, model publishers should describe the expected behaviors and error rates so that users can make minimally informed decisions (mitchell2019model, ). Consistent with the OSS movement, open benchmarks allow for distributed contributions from domain-experts ensuring benchmarks have broad topical coverage. What is more, open contributions distribute the maintenance costs for this evaluation tooling.
Domain Contextualized
The state of the art benchmarking for LLMs evaluates models on tasks that are challenging, yet are very narrow. General reasoning benchmarks evaluate reasoning capabilities across a range of broad tasks such adjudicating boardgames rule questions (kazemi2024, ), complex geometry problems (kazemi2023, ) and causal reasoning (bbeh2025, ). Developers have chosen to drive models toward performance on these narrowly defined tasks with the consequence that the models hallucinate at higher rates in other zero- and few-shot prompts (Zhou2024larger_and_more, ; nyt2025, ).
Simply stated, it is not possible for teams to develop effective benchmark evaluations in areas where they are non-expert. These teams lack knowledge of the key dimensions to consider for a topic outside their expertise as well as the skill to design questions that would evaluate evaluate models on those dimensions. These teams are able to write benchmarks that are difficult in their domain — solvable by the average humans or skilled language model researchers — but they fall short in representing deep knowledge in other fields. The result is that the best current community-generated, contextualized evaluation frameworks focus on reasoning tasks that do not generalize well to novel, domain-specific tasks. When models are developed without human expertise as a part of the training, they hallucinate more frequently when asked for expert-level advice compared to prompts that do not require expertise (Zhou2024larger_and_more, ).
4 Case Study in Hallucination Dataset Development
In order to produce models that are both scientifically valid and useful in real-world settings, we develop a process that facilitates the comparison between expert, human generated text and non-expert, model-generated text.
Question Development
We first develop sixty-four questions related to political procedures and medical diagnoses. For the political domain, we collaborated with the political practitioner to craft both the questions (e.g., “Can I register to vote on election day in California?” and “Do I need ID to vote in Arizona?”) and the corresponding factual, verifiable answers about electoral procedures (see e.g., ncsl2025election, ). In the medical domain, our research team generated questions (e.g., “I have ulcerative colitis. Will mesalamine help control it?” and “I’m experiencing joint pain, morning stiffness, and swollen fingers. Do I have psoriatic arthritis?”) using publicly available resources, and a physician provided plausible diagnoses.
Expert, Ground Truth Answers
We draw on the expertise of one of the authors, a political consultant with a PhD in political science, and a board-certified physician to develop our golden dataset in the domains of politics and medicine. As we discuss shortly, the way we engaged domain experts directly influenced the validity and usefulness of the resulting hallucination metrics.
We take responses provided by domain experts to be ground truth (GT). How GT is conceptualized in each domain reflects real-world complexities and highlights our main argument — domain-specific context is critical for the development of hallucination metrics. In the political domain, questions such as “Who can register to vote in Arizona?” have a single, correct answer that is defined by law. In contrast, in medicine, clinicians are trained to iteratively narrow toward a diagnosis based on patients’ chief complaints, test results, and responses to medical intervention. Often, prior to intervention, there isn’t a single, definitive conclusion. Ideally for performance benchmarking, GT would reflect a single state; however, we believe the GT should reflect the real-world as not-fully-known as it may be.
LLM Generated Answers
To generate LLM responses, we pass the sixty-four prompts to a single-level GPT-3.5 Turbo model via API.
Evaluation Metrics
All responses are individually provided to a local (offline) keep performance indicator evaluation agent. This agent is capable of both reference-based hallucination measurement through traditional NLP metric computations, as well as reference-free hallucination measurement (leveraging a custom fine-tuned version of Llama 3.1 70B). This agent produces a slate of 12 scores per response. We describe these metrics and cite to the grounding source in Appendix A. For reach domain, we compute an aggregate score for each metrics, which is provided in Table 1.
5 Benchmark Results
Table 1: Average Ground Truth (GT) and LLM Hallucination Metrics for Politics & Medicine (Split View)
Domain Source FCD FGR FDF ECS SCD NRR
Politics GT 10.71 1.54 0.00 0.00 0.0000 0.0000
LLM 11.60 0.57 0.00 0.00 0.0441 0.0000
CTC RCR ROUGE-L BLEU METEOR THS
GT 0.0000 0.1486 1.0000 0.9218 0.9998 9.1786
LLM 0.0000 0.8250 0.1568 0.0096 0.1385 11.0307
Domain Source FCD FGR FDF ECS SCD NRR
Medicine GT 14.93 0.00 0.00 0.00 5.2517 0.0000
LLM 4.97 1.78 0.00 0.00 0.7064 0.0000
CTC RCR ROUGE-L BLEU METEOR THS
GT 0.0000 0.3148 1.0000 0.9514 0.9978 14.9292
LLM 0.0000 0.2858 0.0502 0.0030 0.1317 3.1875
Evidence Cited in Responses
The reported metrics suggest that there is a high hallucination potential for political prompts and a relatively lower potential for medical prompts. However, as we will discuss, the LLM’s seemingly superior performance in medicine compared to the ground truth likely reflects professional norms among physicians, who typically communicate diagnoses without citation or explicit justification.
In the political example, the Factual Claim Density (FCD) score is higher for the LLM than for the GT, and the Factual Grounding References (FGR) is lower in the LLM compared to the GT. The interpretation is that more statements could be construed of as real facts (FCD) in the LLM compared to the GT, and that the LLM language appears to be less grounded in evidence (FGR) compared to the GT. In the medical example, the Factual Claim Density (FCD) score is lower for the LLM than for the GT, and the Factual Grounding References (FGR) is higher in the LLM compared to the GT. The interpretation is that fewer statements could be construed of as a real facts (FCD) in the LLM compared to the GT, and that the LLM language may appear to be more grounded in evidence compared to the GT (FGR).
Our assessment is that the physician’s responses, reflective of how he communicates with patients in real life, did not cite sources or explicitly reference underlying medical theory. By comparison, the LLM responses often grounded their claims in recognizable medical knowledge, referencing specific conditions, medications, or treatment protocols. When we followed up with the physician about this, he explained that doctors do not cite their sources in conversations with patients. This approach contrasts with how a political practitioner might cite evidence or policy when addressing questions about procedures (see, e.g. ncsl2025election, ). When we pressed the physician to think about what sources he would cite if he had to, he replied: “med school.” When pushed further, the physician explained that justifications for diagnoses are taught and memorized in medical school, then internalized over years of clinical practice.
This example highlights a deeper issue: expert answers, while grounded in years of training and experience, may appear less grounded-in-truth than the LLM model responses. Without thoughtful prompt design and clear guidelines for what it means for experts to “cite a source,” hallucination benchmarks risk penalizing expert-informed answers while rewarding seemingly evidence-based responses from LLMs.
Structure of Responses
Our findings underscore the importance of aligning evaluation metrics with the specific task at hand. Several metrics — Fictional Disclaimer Frequency (FDF), Explicit Contextualization Score (ECS), Named Reference Ratio (NRR), Contextual Trigger Count (CTC) — scored $0.00.0$ for both GT and LLMs in both domains. This is particularly significant: it implies that neither the GT responses nor the LLM generations employed structured disclaimers, explicit contextualization, or named references. This absence is not likely attributable to model hallucination; instead it is potentially due to the nature of the measurement task itself. The sample prompts and reference responses often consist of straightforward factual statements that lack the stylistic or rhetorical cues these metrics are designed to detect. These bottom-coded hallucination metrics highlight a fundamental limitation in applying conventional fictionality-sensitive metrics to question and answer domains that are firmly grounded in fact rather than fiction.
Realism and Speculative Language
The text generated by the LLM mimics language that reflects greater realism or plausibility, but also includes language that is speculative. In the political example, the LLM responses averaged higher on the Real-world Concept Ratio (RCR) than the GT. This suggests that the LLM invoked more real-world concepts than the GT. However, this may not correspond to accurate grounding. The Speculative Claim Density (SCD) is higher for the LLM responses than for the GT, indicating an increased tendency to include language that hedges or speculates. The LLM produces text that has a higher THS than GT, driven in large part by the relatively elevated FCD and fewer contextual safeguards compared to the GT. The GT responses, in contrast, are shorter, more constrained, and structurally conservative — resulting in a low total hallucination score (THS), but also exhibit limited contextual sophistication (FGR, FDF, and ECS).
The board-certified, human-doctor produced responses that score considerably higher on THS than the model-generated text, principally due to the FCD chasm. A naive interpretation of THS would conclude that the doctor’s responses presented a greater risk of being hallucinated than the model-generated text. Obviously, this cannot be true. In fact, the way that doctors think, reason, and communicate about diagnoses from chief-complaints is simply poorly measured by these hallucination metrics. As doctors professionalize and develop medical expertise they develop heuristics that allow them to produce correct diagnoses quickly. Given limited, and fallible cognition, they prioritize getting answers correct, rather than citing the grounding scientific study that connects a complaint to a diagnoses. And, when faced with chief-complaints, doctors hold in mind a series of possible diagnoses — leading to higher values of SCD — some of which would be ruled out pending the result of a follow-up laboratory tests. For example, when responding to patients’ chief complaints and symptoms, doctors often say, “It’s possible you have some diagnosis. Let’s run some tests.”
Lexical Structure
Consistent with our expectations, LLM outputs appear fluent, but they often deviate in surface form and phrasing from the human-generated reference text. Such findings underscore the limitations of current LLM generations in matching not just factual content, but precise expression and structure. Across both the political and medical domains, metrics that evaluate similarity between generated text and what a human would have written (i.e. ROUGE-L, BLEU and METEOR), GT responses score as near perfect matches to human written responses. In contrast, LLM responses receive considerably lower scores; there is substantial lexical and structural differences between model-generated text and human-written, ground-truth responses.
6 Alternative Positions
Alternative Position 1: Benchmarking is misguided.
Perhaps the negative consequences of hallucination errors are already so limited that the time and carbon spent reducing them could be better directed toward higher-impact problems. Or, perhaps since hallucinations can never be fully eliminated (li2023haluevallargescalehallucinationevaluation, ; banerjee2024llmshallucinateneedlive, ; xu2025, ), continued efforts to minimize them may represent a misallocation of resources. Another possibility is that general-purpose benchmarks will be sufficient to constrain models in specific tasks-settings, thereby minimizing hallucination.
None of these justifications seem tenable. Rising hallucination rates in current-generation, “reasoning” models contradict the notion that this is a moribund problem (Zhou2024larger_and_more, ); leaving a problem unaddressed because it is challenging or has unknown limit properties is inconsistent with the principles in computer science. These models are designed to produce output that is naturally legible to human users who apply them to both increasingly specialized, sophisticated tasks and broad, general tasks. We can be quite certain that the seemingly coherent language produced for general tasks will result in reduced scrutiny of all model-generated text (bender2021dangers, ). The decreased oversight of model output means that when a model hallucinates an incorrect answer, users often will not realize it: either becuase they lack the expertise to assess the response or because they do not devote sufficient cognitive effort to scrutinize it. If language models are used as “discount experts,” then the deleterious consequences of hallucinations are likely to be concentrated among vulnerable, resource-constrained populations. In light of these risks, expert-informed benchmarking remains a critical component of responsible language model development.
Alternative Position 2: Expertise is too expensive.
Perhaps the costs of benchmarking using domain-expert assessments outweigh the benefits of reducing hallucinations. Certainly, expertise is expensive — spot contracts for expert medical, legal and scientific consulting are commonly several hundreds of dollars per hour. What is more, expertise is highly domain specific. Gathering assessment and benchmarking tasks from experts in one domain is unlikely to prove useful even in closely-related, cognate domains. The implication, in the language of the drunkard’s search, is that it might be prohibitively expensive to generate enough lampposts of domain expertise to provide benchmark illumination across all areas of human knowledge.
We acknowledge these costs, but point out that they are minuscule compared to the enormous costs of training current-generation models and paying staff data scientists and engineers. Furthermore, we point out that model training produces models weights that are obsolescing, private goods; model weights are not commonly shared between companies or research teams, and current generation models are inevitably replaced by future generation models. In contrast, domain-expertise is nonrivalrous — the usage by one team does not decease the ability to be used by another team — and it is evergreen. Expertise can be reused to benchmark all generations of models, and is replaced only when the underlying science, law, or art evolves. The implication is that costs to gather expertise can be amortized across a relatively-longer time horizon and if this expertise is shared within an open scientific community and spread across many research groups, it can serve as a public good driving forward all research teams.
Alternative Position 3: Market dynamics will solve this problem.
The creation and provision of open, public data for domain-specific benchmarking is in contrast to closed, competitive market dynamics where firms compete to develop ever more performant language models. Perhaps the closed, competitive landscape is better suited to investing in benchmarks than the open scientific landscape.
While it is true that commercially developed language models are receiving considerable coverage in the press, it is important to remember that the theoretical foundations for these models were laid in open, scientific dialogue (e.g., vaswani2017attention, ). Commercial firms’ advantage has been in marshaling resources to glean text, process through compute to create private model weights, and deliver packaged product to consumers. Benchmarking is quite different from this: there are strong advantages to coordinating rather than competing, benchmarks are non-rivalrous and work best when commonly used within and across teams, and benchmark contains no trappings of being customer-facing and thus does not require a polished front-end. For exactly these reasons it is likely that market dynamics will lead to an under-provision of open domain-contextualized benchmark data, since this data could be used by a competitor to improve their models.
There are some reasons to believe that dynamics that have led to the success of the open-source software paradigm may be applicable to expert benchmarking. Companies’ underlying value arises from applications built on top of core tools, and coordination around a common tool set (or benchmark set) creates the ability for collaboration both within and between firms in a way that supports and enhances development without giving away trade secrets. Like the open source software movement, companies that participate in open benchmarking benefit because they can recruit academic talent that has extensive exposure to tools and frameworks, rather than requiring costly in-house training on proprietary tooling. And, rather than the maintenance of benchmarks being served by a single team at a single organization, this maintenance can be supported by interested domain experts from the global community.
7 Conclusion
In this position paper, we argue that subjecting language models to repeatable, open, and domain-contextualized hallucination benchmarks will advance language model development. Repeatability is essential for internal testing and scientific rigor because tuning models to perform well on narrow external benchmarks often leads to overfitting: strong performance on the benchmark task, but poor generalization to other contexts. Openness in benchmark development is foundational to science and enables users to independently verify claims made by software providers. Moreover, open benchmarks foster collaboration across organizations and between academic and industry groups, facilitate cost-sharing when procuring questions and answers, and invite contributions to be volunteered from domain experts, ensuring benchmarks reflect up-to-date knowledge. Domain-contextualized benchmarks that represent expert thinking are necessary to hallucination metrics remain effective and relevant when models are inevitably applied to increasingly challenging problems. Failure to include expertise risks deploying unhelpful, potential harmful language models.
Appendix A Metric Definitions
We report the Total Hallucination Score (THS) which is a composite measure that includes Factual Claim Density (FCD), Factual Grounding References (FGR), Fictional Disclaimer Frequency (FDF), Explicit Contextualization Score (ECS) [10]. FCD increases the possibility that a statement could contain a hallucination; grounding statements with a fact (FGR), or making a disclaimer that a statement is fictional (FDF & ECS) decreases the possibility that a statement contains a hallucination.
We also report conventional hallucination metrics: Speculative Claim Density (SCD), Named Reference Ratio (NRR), Contextual Trigger Count (CTC), and Real-world Concept Ratio (RCR).
Finally, we report machine summarization measures: Recall-Oriented Understudy for Gisting Evaluation – Longest Common Subsequence (ROUGE-L), Bilingual Evaluation Understudy (BLEU), and Metric for Evaluation of Translation with Explicit Ordering (METEOR). See Table 2 for a detailed description of the reported metrics.
Abbrev. Metric Definition Interpretation
FCD Factual Claim Density Number of claims that appear to be historical, scientific, or verifiable facts per 100 words. A lower FCD suggests fewer statements that could be mistaken for real facts.
FGR Factual Grounding References Number of references to real-world evidence Lower FGR texts are more overtly fictional.
FDF Fictional Disclaimer Frequency Frequency (per 100 tokens) of explicit indications the text is fictional. Higher FDF means texts are framed as not-based-in-fact.
ECS Explicit Contextualization Score Frequency (per 100 tokens) text includes language such as “purely fictional” or “mythical.” Higher ECS indicates stronger distancing from factual framing.
THS Total Hallucination Score Composite: FCD - (FGR + FDF + ECS). Higher THS implies a denser presence of fact-like claims that are not offset by disclaimers or real-world grounding.
SCD Speculative Claim Density Proportion of sentences containing hedging or speculative language (e.g., “might,” “some believe,” “possibly”). Higher SCD may reflect caution in LLMs, but also signals epistemic uncertainty that can blur fact vs. fiction boundaries.
NRR Named Reference Ratio Proportion of sentences that include named entities tied to identifiable sources (e.g., people, organizations, documents). Higher NRR suggests stronger citation patterns; low values often signal decontextualized or imagined references.
CTC Contextual Trigger Count Number of explicit framing cues per 100 tokens (e.g., “according to legend,” “myth says,” “reportedly”). Higher CTC indicates the model is flagging uncertainty or fictionality more clearly.
RCR Real-world Concept Ratio Proportion of noun phrases or terms that refer to concrete, verifiable real-world entities. Higher RCR reflects greater realism or plausibility; low values suggest fictional abstraction or vagueness.
ROUGE-L Recall-Oriented Understudy for Gisting Evaluation Measures the longest common subsequence between generated and reference text. Higher ROUGE-L indicates stronger structural and lexical alignment with the reference.
BLEU Bilingual Evaluation Understudy Precision-oriented score capturing overlap between n-grams in the candidate and reference. High BLEU reflects lexical fidelity to the reference; low BLEU suggests phrasing divergence.
METEOR Metric for Evaluation of Translation with Explicit ORdering Harmonic mean of unigram precision and recall, with penalties for word order and semantic mismatch. More tolerant than BLEU; higher METEOR values indicate better fluency and synonym-aware alignment with the reference.
Table 2: FCD, FGR, FDF, ECS, and the composite THS are drawn from Gosmar & Dahl, 2025 [10]. Additional reference-free metrics (SCD, NRR, CTC, RCR) [33, 55, 60, 45] extend interpretability in contexts lacking ground-truth. ROUGE-L [32], BLEU [44], and METEOR [2] are reference-based metrics standard in natural language evaluation.
Appendix B Performance Benchmarking
Determining how well an LLM performs is a complex problem. Very few industry or research leaders agree on the most important combination of measures or how they should be weighted. Some of the different ways to conceptualize performance include: time-to-train speed, time to n number of tokens predicted for inference or model efficiency measures such as the number of GPUs and memory size required to load and run a model successfully are all different ways of thinking about LLM performance [14]. In the paragraphs that follow, we describe the current state of the art approaches for performance benchmarking.
Training Performance Benchmarking
Due to the enormous compute memory and storage requirements some of the common resource measurements for establishing benchmarks of LLM training performance include: number of GPUs, size of memory, carbon footprint, and cost associated with electricity and personnel. The length of LLM training is primarily governed by the most efficient size of the token corpora to be processed, before overfitting occurs [14]. Interestingly, determining this optimal token size is done by testing training checkpoints on an accuracy metric. Of course, fine-tuning, mixture-of-experts (MoE) and other post-foundational training practices may extend the training process. From a practical sense, benchmarking test suites such as MLPerf provide tools for independent evaluation of LLM training processes[36].
Inference Performance Benchmarking
Some of the same performance constraints from training are applicable to inference — such as model size, memory, and GPU compute requirements. In addition to this, delivering an LLM’s output to a user (or more practically millions of concurrent users), in an acceptable time is both an engineering problem and a product issue. As such most performance benchmarks are highly subjective and are always local to the model architecture and hardware used. In order to achieve acceptable performance, trained models are often quantized and curated to either reduce the number of unused layers in the model and compress the model, or by reducing the precision of model weights [14, 61]. [^5] MLPerf offers well-documented benchmark tools for evaluation of LLM inference[47].
Appendix C Accuracy Benchmarking
Accuracy benchmarking is an actively developing research area [14, 29]. LLM systems vary in size, domain, technique, technology, applied use case and interface; as a result, evaluating the accuracy of model for accuracy is difficult, and no agreed-up measure or set of metrics exits. In the paragraphs that follow, we describe the current state of the art approaches for accuracy benchmarking.
Training Accuracy Benchmarking
Training accuracy depends primarily on the number of tokens, number of model parameters (model size) and number of GPU’s required to successfully complete the training process [14]. Rather than a single training run, frequently, model checkpoints are created after specific stopping criteria are met; then, metric snapshots from those checkpoints are used to determine future tuning and model optimization practices. Training performance is incredibly important to the field of GenAI and training outcomes are central to the in-the-wild performance of LLMs.
Inference Benchmarking
Hodak et al. highlight the emerging use of Recall-Oriented Understudy for Gisting Evaluation (ROGUE) [14] for text summarization task evaluation. ROGUE is a measure of similarity between human generated (gold standard) summarizations and the synthetic model outputs. Machine translation tasks are often evaluated using Bilingual Evaluation Understudy (BLEU) metrics and datasets, which also measures the similarity between reference translations and AI-generated output. In a more general sense, many research and industry applications leverage the 57 curated tasks from the Massive Multitask Language Understanding (MMLU) project as way to get a benchmark for LLM performance on common actions and knowledge [12]. In contrast, Zhou et al. actually introduce five time-of-inference benchmarks (addition, anagram, locality, science and transforms) which evaluate simple numeracy, vocabulary reshuffling, geographic knowledge, low-complexity to high-complexity scientific knowledge and multitask ability [61]. They then categorize those tasks with relationship to a given model’s ability answer correctly or avoid giving a direct answer. They they then implement these metrics and establish a set of benchmarks for each, based on human-judged difficulty of questions (prompts) given to a variety of models.